OpenStudy (anonymous):

complete the square to solve the equation. (check all that apply) x^2+8x-2=18 a. -3 b. -10 c. -2 d. 2

OpenStudy (lalaly):

(x+4)^2-16-2=18 (x+4)^2=36 x+4 = 6 x=2